OpenStudy (abmon98):

express in partial fraction 1/x(a-x)

OpenStudy (abmon98):

-A/x - B/x-a=1/x(a-x) -A(x-a)-B(x)=1 -Ax+Aa-Bx=1 Collect powers -A-B=0 +Aa=1 A=1/a -1/a-B=0 B=-1/a

OpenStudy (abmon98):

so its 1/ax-1/a(x-a) but iam not sure if this is correct or not

OpenStudy (anonymous):

Im curious as to how you arrived at negative signs?

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 1 }{ x(a-x) }\] Using partial fractions this will become \[\frac{ A }{ x }+ \frac{ B }{ (a-x) } = \frac{ 1 }{ x(a-x)}\]

OpenStudy (anonymous):

Now we would cross multiply \[\frac{ A }{ x }+\frac{ B }{ (a-x) } = A(a-x)+B(x)\] Since the bottom terms are both the same we have \[A(a-x) +B(x) = 1\]

OpenStudy (anonymous):

Although it is stated "express in partial fraction" so we may not be asked to solve for the partial fraction yet it is good practice to know know.

myininaya (myininaya):

You can see if it is write by combining the fractions @Abmon98 Test: \[\frac{1}{ax}-\frac{1}{a(x-a)}=\frac{1(x-a)}{ax(x-a)}-\frac{1(x)}{ax(x-a)}=\frac{x-a-x}{ax(x-a)}=\frac{-a}{ax(x-a)}\\ =\frac{-1}{x(x-a)}=\frac{1}{x(a-x)}\] Therefore the decomposition you have works :) Great job.

myininaya (myininaya):

right not write*

OpenStudy (abmon98):

Thank you @myininaya and @Johnbc for your help :D